Circular No. 5823

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993S IN ANONYMOUS GALAXY M. Hamuy, Cerro Tololo Interamerican Observatory, reports that a

spectrogram of SN 1993S (cf. IAUC 5812) obtained on June 26.27 UT with the CTIO 1.5-m telescope telescope shows broad Hα emission with the characteristic P-Cyg profile of type-II supernovae. The spectrum of the underlying galaxy shows narrow Hα emission yielding a redshift z = 0.032.

V1357 CYGNI M. D. Leising, Clemson University; and M. S. Strickman, W. N. John-

son, J. D. Kurfess, and J. E. Grove, Naval Research Laboratory, report for the Compton Observatory/OSSE Team: “OSSE observed V1357 Cyg (Cygnus X-1) during two intervals, May 13-14 and May 31-June 3, during the period when it was reported to be in a low hard x-ray state (IAUC 5813). We confirm that the 45- to 140-keV flux was low during these periods: 0.08 and 0.06 photon cm^{−2} s^{−1}, respectively. We find no evidence for MeV emission in either observation, with 99-percent confidence upper limits on the integral flux from 0.6 to 1.5 MeV of 2.8 and 1.9 × 10^{−3} photon cm^{−2} s^{−1}. These limits are factors of 4 and 6, respectively, below the flux observed by Ling et al. (1987, Ap.J. 321, L117), when the hard x-ray flux was similarly low.”

RX PUPPIS R. J. Ivison and E. R. Seaquist, University of Toronto, report: “Con-

tinuum measurements obtained with the James Clerk Maxwell Telescope and the Very Large Array during 1992 October and 1993 May, respectively, suggest that the symbiotic binary RX Pup has returned to a state of low excitation after more than a decade of activity. The centimeter spectral index is –0.1 ± 0.1, with a flux density of 13.3 ± 0.7 mJy at 3.6 cm. Spectra obtained with the International Ultraviolet Explorer and the Anglo-Australian Telescope during 1993 April and May confirm that high-excitation emission lines seen during the 1980s have disappeared, leaving only the Balmer series, He I, Fe II, O I, and weak, residual [O III] in emission. The IUE Fine Error Sensor magnitude remains at 11.5, but a weak, red continuum has replaced the emission lines of N IV], C IV, and He II usually seen in the SWP and LWP cameras.”

1993 June 28 Daniel W. E. Green